# Grounded-VideoLLM: Sharpening Fine-grained Temporal Grounding in Video Large Language Models

### v1

@@ 3.2 Unified Temporal Tokens
Given a text span depicting a particular video clip and its associated timestamps, we employ a relative time representation that converts continuous timestamps into a sequence of discrete temporal tokens. For a video $\mathcal{V}$ with a duration of $L$ seconds, we evenly divide $\mathcal{V}$ into $M$ equal-length chunks, and then define $M+1$ anchor points (ranging from <0> to <M>) across $\mathcal{V}$, representing relative temporal positions. Each anchor point corresponds to a specific timestamp of $\mathcal{V}$ and is encoded as a temporal token. For instance, <0> denotes the very start of $\mathcal{V}$ while <M> indicates the end. These $M+1$ tokens are added to the LLM’s vocabulary to enable unified modeling alongside text. A specific continuous timestamp $\tau$ can be easily converted to a temporal token <t> and vice verse:
@@
While this may introduce minor quantization errors, these can be minimized by selecting an appropriate $M$. We then organize the text span and its corresponding temporal tokens in a unified format. Both text tokens and temporal tokens are mapped to embeddings through the extended word embedding layer of LLM. For example, one input representation is as follows:
<s><video>$\textbf{F}_{Vid}$</video> <grounded> From <0> to <6>, a baby is crying. From <7> to <16>, a man is coming and picking up the baby. From <20> to <25>, the baby is eating an apple. From <27> to <35>, the baby is smiling happily.</s>
where <s> and </s> indicate start- and end-of-sequence, <video> and </video> represent the beginning and end of encoded video representations. <grounded> is a special token to tell the model should output the grounded timestamps. This strategy avoids the need to tokenize and process numerical values, which has been identified as a limitation of LLMs (Schwartz et al., 2024), and it greatly simplifies the representation of timestamps within the unified embedding space of LLMs. Consequently, text and timestamps can be jointly decoded as a single sequence, following the objective outlined in Eq. (2).
4 Progressive Training
Different from previous methods (Zhang et al., 2023b; Lin et al., 2023) that train models from scratch using mixed image and video datasets, we start with a pre-trained image-based MLLM (Microsoft, 2024) and progressively enhance its fine-grained temporal grounding capabilities. This strategy can be applied to any off-the-shelf MLLM and is more efficient. Table 1 enumerates the datasets used at different training stages, and Table 9 lists the prompts for different tasks.
Stage-1: Video-Caption Alignment. Feature alignment is widely used to improve training efficiency (Liu et al., 2023). In this stage, we leverage approximately 1.28 million video-text pairs sampled from diverse sources (Wang et al., 2024a; Bain et al., 2021; Chen et al., 2024b) to align the video encoder’s features with the MLLM. This alignment allows the MLLM, which was pre-trained solely on images, to gain a foundational understanding of videos. Since 2D down-sampling has been applied to the visual features, only the two projectors $f(\cdot)$ and $g(\cdot)$ are set to be trainable, while the video encoder, image encoder, and LLM remain frozen. As this stage does not involve any video temporal grounding tasks, the temporal tokens described in Sec.3.2 are not yet incorporated.
Stage-2: Temporal Token Alignment. While video-caption alignment effectively connects videos and the MLLM at a coarse semantic level, a gap persists between this alignment and fine-grained temporal grounding. To address this, we introduce the temporal tokens described in Sec.3.2 and continue pre-training the model on a diverse range of grounding datasets (Huang et al., 2024a; Qian et al., 2024a; Wang et al., 2024c), focusing on tasks such as Temporal Sentence Grounding, Dense Video Captioning, and Temporal Referring, which enables the model to refer to and localize temporal information effectively. Since new tokens are introduced, the trainable parameters in this stage include the two projectors, $f(\cdot)$ and $g(\cdot)$, the word embedding matrix, and the final classifier head of the LLM. This step enhances the model’s ability to comprehend multiple events and aligns the temporal tokens with both the video timelines and the LLM’s semantic space.
Stage-3: Multi-Task Instruction Tuning. Following the initial two stages of pre-training, the model has developed a basic understanding of video content and the ability to refer to and locate specific timestamps. In this stage, we will further enhance the model’s fine-grained temporal grounding while improving its responsiveness to diverse user instructions. To achieve this, we gather two types of datasets: (1) We compile a wide range of public datasets for video temporal grounding tasks, similar to Time-IT (Ren et al., 2024) and VTG-IT (Guo et al., 2024), which include tasks of dense video captioning, temporal sentence grounding, and grounded VideoQA. (2) We incorporate a selection of instructional video-to-text datasets from VideoChat2 (Li et al., 2024), which feature tasks such as conversations, classification, question answering, and captioning. Additionally, we include ShareGPT-4Video (Chen et al., 2024a) to further enhance the model’s ability to generate detailed video captions. By utilizing these diverse datasets, which encompass both temporal grounding and video instruction tasks, Grounded-VideoLLM excels in temporal referring, localization, reasoning, and general comprehension of video content. In this stage, the trainable parameters remain the same as in Stage 2, with the addition of LoRA parameters (Hu et al., 2022) for the LLM.
@@ 4 Progressive Training
Multi-Task Instruction Tuning Grounded Conversation 442K ANet-RTL, Moment-10M
@@
Table 1: Datasets used at three training stages. Tasks with gray background consist of datasets regarding temporal grounding.
@@
5 Grounded VideoQA Dataset Generation
Grounded VideoQA requires the model to not only answer questions but also identify relevant video moments that support predicted answers, thereby demonstrating the model’s temporal reasoning abilities. The NExT-GQA dataset (Xiao et al., 2024) was manually developed by extending NExT-QA (Xiao et al., 2021) with temporal labels for start and end timestamps. However, annotating these temporal labels is labor-intensive and time-consuming, which has limited NExT-GQA only to QA pairs for the validation and test sets. To create a scalable training dataset, we utilized OpenAI GPT-4 (Achiam et al., 2023) to assist in constructing training sets for the grounded VideoQA task. These sets were built on public datasets that already contain temporal labels, such as ActivityNet-Caption (Caba Heilbron et al., 2015) and QVHighlights (Lei et al., 2021). We framed the task as a multiple-choice problem using a two-round conversational format, as depicted in Figure 3.
6 Experiments
6.1 Experiment Setting
Implementation Details. We select the Phi3.5-Vision-Instruct-3.8B (Microsoft, 2024) as the base MLLM for our Grounded-VideoLLM. For temporal stream encoding, we adopt InternVideo2-1B (Wang et al., 2024b) as the video encoder. Each video is sampled as a sequence of $T=96$ frames, which are evenly divided into $K=12$ segments. We set the pooling size $\sigma_{S}=2$ for the spatial stream while $\sigma_{T}=4$ for the temporal stream respectively, which results in $N_{S}=144$ tokens per frame for the spatial stream while $N_{T}=16$ tokens per frame for the temporal stream. Moreover, we introduce $M=300$ temporal tokens into the LLM’s vocabulary for timestamp representation. More implementation details can be found in Appendix A.1.
Tasks and Benchmarks. To thoroughly evaluate Grounded-VideoLLM in fine-grained temporal grounding, we assess it across three video temporal grounding tasks: Temporal Sentence Grounding, Dense Video Captioning, and Grounded VideoQA, utilizing datasets such as Charades-STA (Gao et al., 2017), ActivityNet-Captions (Caba Heilbron et al., 2015), and NExT-GQA (Xiao et al., 2024). We also show its reasoning capability by the task of Open-Ended VideoQA with datasets including MSVD-QA, MSRVTT-QA (Xu et al., 2017), and ActicityNet-QA (Yu et al., 2019). Additionally, to evaluate the model’s general video understanding capabilities, we benchmark Grounded-VideoLLM against existing models using modern Video-LLM benchmarks including VCG-Bench (Maaz et al., 2024b) and MVBench (Li et al., 2024).
@@ 6.1 Experiment Setting
Grounded-VideoLLM 4B 54.2 36.4 19.7 36.8 46.2 30.3 19.0 36.1 6.0 6.8
Table 2: Zero-shot results on temporal sentence grounding and dense video captioning tasks.
Model Acc@GQA mIoP IoP@0.3 IoP@0.5 mIoU IoU@0.3 IoU@0.5
VIOLETv2 (Fu et al., 2023) 12.8 23.6 25.1 23.3 3.1 4.3 1.3
Temp[CLIP] NG+ (Xiao et al., 2024) 16.0 25.7 31.4 25.5 12.1 17.5 8.9
SeViLA (Yu et al., 2023) 16.6 29.5 34.7 22.9 21.7 29.2 13.8
HawkEye (Wang et al., 2024c) - - - - 25.7 37.0 19.5
LangRepo (Kahatapitiya et al., 2024) 17.1 31.3 - 28.7 18.5 - 12.2
FrozenBiLM NG+ (Yang et al., 2022) 17.5 24.2 28.5 23.7 9.6 13.5 6.1
VideoStreaming (Qian et al., 2024b) 17.8 32.2 - 31.0 19.3 - 13.3
LLoVi (Zhang et al., 2023a) 24.3 37.3 - 36.9 20.0 - 15.3
Grounded-VideoLLM 26.7 34.5 42.6 34.4 21.1 30.2 18.0
@@
[…]
@@ 6.2 Main Results
Temporal Sentence Grounding requires the model to identify the precise time interval corresponding to a given query sentence. As shown in Table 2, Grounded-VideoLLM achieves performance on “mIoU” with 36.8 for the Charades-STA (Gao et al., 2017) and 36.1 for ActivityNet-Grounding (Caba Heilbron et al., 2015) respectively, surpassing previous SoTA end-to-end Video-LLMs, i.e., HawkEye (Wang et al., 2024c), by a significant margin (+3.4). It is worth mentioned that the promising performance of “mIoU” are largely attributed to the signigicant gains in terms of ”R@0.7” compared with other thresholds, demonstrating that Grounded-VideoLLM is more advanced in localizing specific moments within videos with finer granularity.
Dense Video Captioning involves generating descriptions for all events in a video, along with their corresponding start and end timestamps. We evaluated Grounded-VideoLLM on the ActivityNet-Captions (Caba Heilbron et al., 2015), and the results in Table 2 show that our method achieves the highest SODA_c score (6.0), which demonstrates that, thanks to the Temporal Token Alignment training stage, Grounded-VideoLLM is highly effective in identifying the multi-event structure of the video and capturing complete storylines. The highest METEOR score (6.8) also indicates that Grounded-VideoLLM provides more detailed event descriptions compared with other Video-LLMs.
Model MSVD-QA MSRVTT-QA ANet-QA VCG-Bench
Acc. Score Acc. Score Acc. Score CI DO CU TU CO Avg.
@@
Video-LLaMA (Zhang et al., 2023b) 51.6 2.5 29.6 1.8 12.4 1.1 1.96 2.18 2.16 1.82 1.79 1.98
Video-ChatGPT (Maaz et al., 2024b) 64.9 3.3 49.3 2.8 35.2 2.7 2.50 2.57 2.69 2.16 2.20 2.42
Vista-LLaMA (Ma et al., 2024) 65.3 3.6 60.5 3.3 48.3 3.3 2.44 2.64 3.18 2.26 2.31 2.57
MovieChat (Song et al., 2024) 75.2 3.8 52.7 2.6 45.7 3.4 2.76 2.93 3.01 2.24 2.42 2.67
LongVLM (Weng et al., 2024) 70.0 3.8 59.8 3.3 47.6 3.3 2.76 2.86 3.34 2.39 3.11 2.89
VideoChat2 (Li et al., 2024) 70.0 3.9 54.1 3.3 49.1 3.3 3.02 2.88 3.51 2.66 2.81 2.98
Chat-UniVi (Jin et al., 2024) 65.0 3.6 54.6 3.1 45.8 3.2 2.89 2.91 3.46 2.89 2.81 2.99
P-LLaVA-7B (Xu et al., 2024a) 76.6 4.1 62.0 3.5 56.3 3.5 3.21 2.86 3.62 2.33 2.93 3.12
ST-LLM (Liu et al., 2024a) 74.6 3.9 63.2 3.4 50.9 3.3 3.23 3.05 3.74 2.93 2.81 3.15
VideoGPT+ (Maaz et al., 2024a) - - - - - - 3.27 3.18 3.74 2.83 3.39 3.28
@@
Momentor (Qian et al., 2024a) 68.9 3.6 55.6 3.0 40.8 3.2 - - - - - -
VTimeLLM (Huang et al., 2024a) - - - - - - 2.78 3.10 3.40 2.49 2.47 2.85
LITA (Huang et al., 2024b) - - - - - - 2.94 2.98 3.43 2.68 3.19 3.04
Grounded-VideoLLM 76.3 4.1 60.3 3.6 56.8 3.5 3.34 2.94 3.66 3.12 3.14 3.24

### v2

@@ 3.2 Unified Temporal Tokens
Given a text depicting a video and its associated timestamps, we employ a relative time representation that converts continuous timestamps into a sequence of discrete tokens. For a video $\mathcal{V}$ with a duration of $L$ seconds, we evenly divide $\mathcal{V}$ into $M$ equal-length chunks, and then define $M+1$ anchor points (ranging from <0> to <M>) across $\mathcal{V}$, denoting relative temporal positions. Each anchor point corresponds to a specific timestamp and is encoded as a temporal token. For instance, <0> denotes the very start of $\mathcal{V}$ while <M> indicates the end. These $M+1$ tokens are added to the LLM’s vocabulary to enable unified modeling alongside text. A specific continuous timestamp $\tau$ can be easily converted to a temporal token <t> and vice verse:
@@
While this may introduce minor quantization errors, these can be minimized by selecting an appropriate $M$ or an interpolation expansion. We then organize the text span and its corresponding temporal tokens in a unified format. Both text tokens and temporal tokens are mapped to embeddings through the extended word embedding layer of LLM. For example, one input representation is as follows:
$\textbf{F}_{Vid}</video$> <grounded> From <0> to <6>, a baby is crying. From <7> to <16>, a man is coming and picking up the baby. From <20> to <25>, the baby is eating an apple.
where <video> and </video> represent the beginning and end of encoded video representations. <grounded> is a special token to tell the model should output the grounded timestamps.
@@
Multi-Task Instruction Tuning Grounded Conversation 442K RTL, Moment-10M
@@
Table 1: Datasets at three training stages. Tasks with gray consist of datasets regarding temporal grounding.
This strategy avoids the need to tokenize and process numerical values, which has been identified as a limitation of LLMs (Schwartz et al., 2024). Notably, special tokens in LLMs are widely used in various domains. For example, Pix2Seq (Chen et al., 2021) leverages special tokens to represent spatial grounding in images, Open-VLA (Kim et al., 2024) and RT-2 (Brohan et al., 2023) utilize them for encoding robot action spaces, Yo’LLaVA (Nguyen et al., 2024) use special tokens to refer to personalized subjects, and, most relevant to our work, Vid2Seq employs special tokens for temporal grounding in videos. However, the aforementioned works lack effective strategies to align these tokens with both LLM’s semantic meanings and specific functionalities. For example, Vid2Seq (Yang et al., 2023) simply appends temporal tokens as a prefix to the caption and trains the entire T5 model from scratch using noised transcribed speech, which disrupts the language model’s original semantic embedding. In contrast, we introduce a temporal token alignment training stage in Sec. 4 to mitigate this issue. Instead of training the entire model, we update only the word embeddings of temporal tokens and the final logit head, with carefully curated grounding-specific datasets. This ensures that temporal tokens are aligned with both the video timeline and the LLM’s semantic space, and timestamps and text can be jointly decoded as a single sequence while maintaining the general video understanding ability.
@@
4 Progressive Training
Different from previous methods (Zhang et al., 2023b; Lin et al., 2023) that train models from scratch using mixed image and video datasets, we start with a pre-trained image-based MLLM (Microsoft, 2024) and progressively enhance its fine-grained temporal grounding capabilities. This strategy can be applied to any off-the-shelf MLLM and is more efficient. Table 1 enumerates the datasets used at different stages.
Stage-1: Video-Caption Alignment. We leverage approximately 1.28 million video-text pairs sampled from diverse sources (Wang et al., 2024a; Bain et al., 2021; Chen et al., 2024b) to align video encoder’s features with the MLLM. This alignment allows the MLLM, which was pre-trained solely on images, to gain a foundational understanding of videos. Since 2D down-sampling has been applied to the visual features, only the two projectors $f(\cdot)$ and $g(\cdot)$ are trainable, while the video encoder, image encoder, and LLM remain frozen. As this stage does not involve any temporal grounding tasks, the temporal tokens in Sec.3.2 are not yet incorporated.
Stage-2: Temporal Token Alignment. While Video-Caption Alignment connects videos and the MLLM at a coarse level, a gap persists between this alignment and fine-grained temporal grounding. To address this, we introduce the temporal tokens described in Sec.3.2 and continue pre-training the model on a diverse range of grounding datasets (Huang et al., 2024a; Qian et al., 2024a; Wang et al., 2024c), focusing on tasks such as Temporal Sentence Grounding, Dense Video Captioning, and Temporal Referring, which enables the model to refer to and localize temporal information effectively. Since new tokens are introduced, we additionally make the word embedding matrix and the final classifier head of the LLM trainable. This stage enhances the model’s ability to comprehend multiple events and aligns the temporal tokens with both the video timelines and the LLM’s semantic space as shown in Figure 4 and Table 5, distinguishing us from previous works Huang et al. (2024b); Yang et al. (2023); Qian et al. (2024a).
Stage-3: Multi-Task Instruction Tuning. Following the initial two stages, the model has developed a basic understanding of video content and can locate specific timestamps. In this stage, we further enhance the model’s fine-grained temporal grounding while improving its responsiveness to diverse instructions. To achieve this, we gather two types of datasets: (1) We compile a wide range of datasets for temporal grounding tasks, similar to Time-IT (Ren et al., 2024) and VTG-IT (Guo et al., 2024), which include tasks of dense video captioning (remove ActivityNet (Caba Heilbron et al., 2015)), temporal sentence grounding, and grounded VideoQA. (2) We incorporate a subset of instructional datasets from VideoChat2 (Li et al., 2024) and ShareGPT-4Video (Chen et al., 2024a) to further enhance the model’s ability to generate detailed video captions. By utilizing these diverse datasets, which encompass both temporal grounding and video instruction tasks, Grounded-VideoLLM excels in temporal referring, localization, reasoning, and general comprehension of video content. In this stage, the trainable parameters remain the same as in Stage 2, with the addition of LoRA parameters (Hu et al., 2022) for the LLM.
@@ 4 Progressive Training
Grounded-VideoLLM (Vicuna) 7B 51.8 34.3 18.3 34.7 43.9 29.1 18.3 34.5 6.2 6.4
Grounded-VideoLLM (Phi3.5) 4B 54.2 36.4 19.7 36.8 46.2 30.3 19.0 36.1 6.0 6.8
Table 2: Zero shot results on temporal sentence grounding and dense video captioning tasks.
Model Acc@GQA mIoP IoP@0.5 mIoU IoU@0.5
VIOLETv2 (Fu et al., 2023) 12.8 23.6 23.3 3.1 1.3
Temp[CLIP] NG+ (Xiao et al., 2024) 16.0 25.7 25.5 12.1 8.9
SeViLA (Yu et al., 2023) 16.6 29.5 22.9 21.7 13.8
LangRepo (Kahatapitiya et al., 2024) 17.1 31.3 28.7 18.5 12.2
FrozenBiLM NG+ (Yang et al., 2022) 17.5 24.2 23.7 9.6 6.1
VideoStreaming (Qian et al., 2024b) 17.8 32.2 31.0 19.3 13.3
LLoVi (Zhang et al., 2023a) 24.3 37.3 36.9 20.0 15.3
Grounded-VideoLLM (Vicuna) 24.0 32.2 31.2 20.8 16.9
Grounded-VideoLLM (Phi3.5) 26.7 34.5 34.4 21.1 18.0
@@
5 Grounded VideoQA Dataset Generation
Grounded VideoQA requires the model to not only answer questions but identify timestamps that support predicted answers, demonstrating the temporal reasoning abilities. NExT-GQA (Xiao et al., 2024) was manually developed by extending NExT-QA (Xiao et al., 2021) with temporal labels for start and end timestamps. However, annotating these labels is labor-intensive and time-consuming, limiting NExT-GQA only to QA pairs for the validation and test sets. To create a scalable training dataset, as depicted in Figure 3, we utilized OpenAI GPT-4 (Achiam et al., 2023) to assist in constructing training sets for the grounded VideoQA task. These sets were built on public datasets that already contain temporal labels, such as QVHighlights (Lei et al., 2021). We framed the task as a multiple-choice VideoQA using a two-round conversational format as depicted in Appendix A.3.
6 Experiments
Implementation Details. We select Phi3.5-V-Instruct-3.8B (Microsoft, 2024) as the base MLLM of Grounded-VideoLLM. We also build another Grounded-VideoLLM based on LLaVA-1.5-7B (Liu et al., 2024a) using Vicuna-1.5 (Chiang et al., 2023) as the LLM for fair comparison. For temporal stream, we adopt InternVideo2-1B (Wang et al., 2024b) as the video encoder. Each video is sampled as a sequence of $T=96$ frames, which are evenly divided into $K=12$ segments. We set the pooling size $\sigma_{S}=2$ for the spatial stream ($N_{S}=144$ tokens per frame) while $\sigma_{T}=4$ ($N_{T}=16$ tokens per frame) for the temporal stream. Moreover, we introduce $M=300$ temporal tokens into the LLM’s vocabulary for timestamp representation. We use the Phi3.5 version for ablations in Section 6.2. More details are in Appendix A.1.
Tasks and Benchmarks. We assess Grounded-VideoLLM across three video temporal grounding tasks: Temporal Sentence Grounding, Dense Video Captioning, and Grounded VideoQA, utilizing benchmarks such as Charades-STA (Gao et al., 2017), ActivityNet-Captions (Caba Heilbron et al., 2015), and NExT-GQA (Xiao et al., 2024). We also show its reasoning capability by the task of Open-Ended VideoQA with benchmarks including MSVD-QA, MSRVTT-QA (Xu et al., 2017), and ActicityNet-QA (Yu et al., 2019). Additionally, to evaluate the model’s general video understanding capabilities, we benchmark Grounded-VideoLLM against existing models using VCG-Bench (Maaz et al., 2024b) and MVBench (Li et al., 2024). The evaluation details are in Appendix A.4.
@@ 6.1 Main Results
Temporal Sentence Grounding. As shown in Table 2, Grounded-VideoLLM (Phi3.5) achieves "mIoU" scores of 36.8 on Charades-STA and 36.1 on ANet-Grounding, respectively. This performance surpasses previous SoTA end-to-end Video-LLMs, such as HawkEye (Wang et al., 2024c), by a substantial margin of +3.4. It is worth emphasizing that this promising "mIoU" performance is largely attributed to significant gains in the "R@0.7" metric compared to other thresholds, demonstrating that Grounded-VideoLLM is more advanced in localizing specific moments with finer granularity. Interestingly, although Grounded-VideoLLM (Vicuna) may be larger than the Phi3.5 version in terms of model parameters, its overall performance is slightly lower. This is because the base Vicuna model is inherently weaker than Phi3.5. Nevertheless, when using the same LLM as the base model, Grounded-VideoLLM (Vicuna) still outperforms other models like TimeChat and VTimeLLM.
Dense Video Captioning. We evaluated Grounded-VideoLLM on the ANet-Captions, and the results in Table 2 show that Grounded-VideoLLM (Phi3.5) achieves the highest SODA_c score of 6.0, which demonstrates that, thanks to the Temporal Token Alignment training stage, Grounded-VideoLLM is highly effective in identifying the multi-event structure of the video and capturing complete storylines. The highest METEOR score (6.8) also indicates that Grounded-VideoLLM provides more detailed event descriptions compared with other Video-LLMs.
NExT-GQA (Xiao et al., 2024) requires the model to both correctly answer questions and provide timestamps that support the answers, highlighting the temporal reasoning capability. According to Table 3, Grounded-VideoLLM achieves the highest Acc@GQA (26.7, +2.4) and delivers comparable IoU and IoP scores to models such as SeViLA (Yu et al., 2023) and LLoVi (Zhang et al., 2023a), which use specialized grounding modules or rely on proprietary large language models (Achiam et al., 2023). The highest Acc@GQA score further demonstrates Grounded-VideoLLM’s capability in both fine-grained temporal grounding and high-level reasoning.
General Video-LLM Benchmarks. While Grounded-VideoLLM excels in fine-grained temporal grounding, we should ensure that it maintains performance in general video understanding. Therefore, we conducted a comprehensive evaluation using VCG-Bench (Maaz et al., 2024b) and MVBench (Li et al., 2024). As shown in Table 4, Grounded-VideoLLM achieves promising results in VCG-Bench, with an average score of 3.24, outperforming other Video-LLMs with temporal grounding capabilities (e.g., LITA, VTimeLLM). Notably, Grounded-VideoLLM surpasses all other Video-LLMs on the TU (Temporal Understanding) task (see Appendix Table 8), with a score of 3.12 (+7%), demonstrating its superior temporal understanding, which can be attributed to the two-stream architecture that can capture motion dynamics. For MVBench which provides 4,000 QA pairs spanning a wide range of scenes categorized into 20 fine-grained tasks, the results, presented in Table 4, show that Grounded-VideoLLM achieves an average score of 60.0, surpassing other Video-LLMs.
Model MSVD-QA MSRVTT-QA ANet-QA VCG-Bench MVBench
Acc. Score Acc. Score Acc. Score Avg. Avg.
@@
Video-LLaMA (Zhang et al., 2023b) 51.6 2.5 29.6 1.8 12.4 1.1 1.98 34.1
Video-ChatGPT (Maaz et al., 2024b) 64.9 3.3 49.3 2.8 35.2 2.7 2.42 32.7
Video-LLaVA (Lin et al., 2023) 70.7 3.9 59.2 3.5 45.3 3.3 - 43.0
Vista-LLaMA (Ma et al., 2024) 65.3 3.6 60.5 3.3 48.3 3.3 2.57 -
MovieChat (Song et al., 2024) 75.2 3.8 52.7 2.6 45.7 3.4 2.67 -
LongVLM (Weng et al., 2024) 70.0 3.8 59.8 3.3 47.6 3.3 2.89 -
VideoChat2 (Li et al., 2024) 70.0 3.9 54.1 3.3 49.1 3.3 2.98 51.1
Chat-UniVi (Jin et al., 2024) 65.0 3.6 54.6 3.1 45.8 3.2 2.99 -
P-LLaVA-7B (Xu et al., 2024a) 76.6 4.1 62.0 3.5 56.3 3.5 3.12 46.6
ST-LLM (Liu et al., 2024b) 74.6 3.9 63.2 3.4 50.9 3.3 3.15 54.9
VideoGPT+ (Maaz et al., 2024a) - - - - - - 3.28 58.7
@@
TimeChat (Huang et al., 2024b) - - - - - - - 38.5
Momentor (Qian et al., 2024a) 68.9 3.6 55.6 3.0 40.8 3.2 -
VTimeLLM (Huang et al., 2024a) - - - - - - 2.85
LITA (Huang et al., 2024b) - - - - - - 3.04
Grounded-VideoLLM (Vicuna) 74.7 3.9 61.9 3.6 55.7 3.4 3.26 58.1
Grounded-VideoLLM (Phi3.5) 76.3 4.1 60.3 3.6 56.8 3.5 3.24 60.0
Table 4: Results on VideoQA, VCG-Bench and MVBench. Refer to Appendix Table 9, 8 for details.
Figure 4: Attention weights of the LLM when generating the temporal tokens and 3D-PCA of embeddings.
6.2 In-Depth Analysis
Two-stream Encoding. We conduct ablations to our two-stream encoding. Specifically, we set three variants by removing the temporal stream or spatial stream respectively: (1) w/o temporal-stream (dense) feeds $T=96$ frames with a pooling size $\sigma_{S}=4$ (36 tokens per frame). (2) w/o temporal-stream (sparse) feeds $T=24$ frames with a pooling size $\sigma_{S}=2$ (144 tokens per frame). (3) w/o spatial-stream feeds $T=14$ frames without pooling (256 tokens per frame). All these variants have a close number of video tokens compared to our two-stream encoding ($12\times(144+\frac{96}{12}\times 16)=3264$ tokens for $\textbf{F}_{Vid}\in\mathbb{R}^{K\cdot(N_{S}+\frac{T}{K}\cdot N_{T})\times D}$) for a fair comparison. Table 5 shows that, interestingly, the dense frame variant performs slightly better in temporal grounding tasks while worse in general benchmarks than the sparse frame variant. This can be attributed to that the videos in MVBench are much shorter and emphasize spatial details more. Our two-stream architecture strikes a balance by capturing dense motion dynamics while maintaining essential appearance details.
Temporal Tokens. We conducted ablations on the temporal tokens to study how they will affect the grounding performance. The models in Table 6 were trained using only the first two stages and evaluated on grounding tasks. Specifically, ’w/o temporal tokens’ refers to directly using plain text to represent absolute timestamps. The results in Table 6 show that while the performance of plain text and 100 temporal tokens is comparable, both are outperformed by 300 tokens. Furthermore, the results reveal a consistent improvement in performance with an increasing number of temporal tokens, especially for longer videos (ANet), highlighting the benefit of finer-grained time representations.
Model # of video C-STA ANet-G ANet-Cap MVBench
tokens mIoU mIoU SODA_c Avg.
Grounded-VideoLLM 3264 36.8 36.1 6.0 60.0
w/o temporal-stream (sparse) 3456 30.4 28.0 4.9 58.5
w/o temporal-stream (dense) 3456 34.3 29.2 5.4 53.2
w/o spatial-stream 3584 33.5 28.7 5.5 57.7
w/o temporal token alignment 3264 27.5 23.1 4.7 58.9
Table 5: Impact of two-stream and token alignment.
Setting C-STA (mIoU) ANet-G (mIoU) ANet-Cap (SODA_c)
w/o temporal tokens 32.3 29.6 5.6
w/ 100 temporal tokens 32.2 29.1 5.2
w/ 200 temporal tokens 32.9 30.1 5.5
w/ 300 temporal tokens 33.8 33.1 5.7
Table 6: Impact of temporal tokens.
Alignment Training Stage. We investigate temporal tokens’ role by ablating the 2nd training stage of Temporal Token Alignment. Quantitative results in Table 5 reveal a performance drop across all tasks, particularly in temporal sentence grounding. Upon analyzing the outputs, we found that the model often produces time intervals spanning nearly the entire video (e.g., from <0> to <300>), neglecting the alignment between specific moments and temporal tokens. Qualitatively, we visualize the attention weights of the LLM to demonstrate how temporal tokens attend to corresponding video moments. Details of generating visualizations are provided in Appendix A.5. As shown in Figure 4 (a), when generating the temporal token, e.g. <241> or <271>, the attention weights are higher and more focused on their corresponding video moments. Conversely, in Figure 4 (b), when the model is trained without the alignment stage, the attention weights of temporal tokens become significantly dispersed across irrelevant moments. This illustrates the necessity of our multi-stage training strategy for temporal alignment. We also visualize the embedding distribution of temporal tokens with PCA in Figure 4 (c), revealing that temporal tokens with similar indices tend to cluster together, exhibiting a continuous transition from tokens with smaller indices to larger ones.
Model NExT-GQA
Acc@GQA mIoP mIoU
Grounded-VideoLLM 26.7 34.5 21.1
Table 7: Impact of grounded VideoQA dataset.
Grounded VideoQA Dataset. We validate the role of our constructed grounded-VideoQA by removing it from the stage-3 data. Since the model without training on our dataset usually generates free-form texts when asked to output the timestamps supporting the answer, we reformulate it as a temporal sentence grounding task, where we combine the predicted answer and question into a single sentence and ask the model to localize its timestamps. Table 7 suggests that there is a significant performance decrease with regard to Acc@GQA ($\downarrow$ 8.6), mIoP ($\downarrow$ 12.3), and mIoU ($\downarrow$ 8.2), from which we can conclude that our Grounded VideoQA dataset is essential for the model’s temporal reasoning capability.
7 Conclusion
We present Grounded-VideoLLM, a novel Video-LLM that incorporates a two-stream encoding for temporal modeling, along with the temporal tokens for timestamp representation. We employ a multi-stage training scheme, starting with an image-based MLLM and progressively equipping it with fine-grained temporal grounding capabilities. Additionally, we curate a grounded-VideoQA dataset to further enhance the model’s temporal reasoning ability. Extensive experiments demonstrate that Grounded-VideoLLM not only excels in video temporal grounding tasks but also performs strongly on general video understanding benchmarks.
8 Limitations
While Grounded-VideoLLM demonstrates superiority in handling fine-grained temporal grounding, but it still has some inherent limitations for future works. (1) Timestamp Quantization Error: Although discrete temporal tokens are introduced to represent timestamps and accuracy is improved by increasing the number of tokens, minor quantization errors may still be introduced when converting continuous time into discrete tokens. (2) Computational Resource Requirements: The training and inference processes of the model, especially the parts involving two-stream encoding and large-scale language models, may require more computational resources than single-stream.
Appendix A Appendix
Model VCG-Bench
CI DO CU TU CO Avg.
Video-LLaMA (Zhang et al., 2023b) 1.96 2.18 2.16 1.82 1.79 1.98
Video-ChatGPT (Maaz et al., 2024b) 2.50 2.57 2.69 2.16 2.20 2.42
Vista-LLaMA (Ma et al., 2024) 2.44 2.64 3.18 2.26 2.31 2.57
MovieChat (Song et al., 2024) 2.76 2.93 3.01 2.24 2.42 2.67
LongVLM (Weng et al., 2024) 2.76 2.86 3.34 2.39 3.11 2.89
VideoChat2 (Li et al., 2024) 3.02 2.88 3.51 2.66 2.81 2.98
Chat-UniVi (Jin et al., 2024) 2.89 2.91 3.46 2.89 2.81 2.99
P-LLaVA-7B (Xu et al., 2024a) 3.21 2.86 3.62 2.33 2.93 3.12
ST-LLM (Liu et al., 2024b) 3.23 3.05 3.74 2.93 2.81 3.15
VideoGPT+ (Maaz et al., 2024a) 3.27 3.18 3.74 2.83 3.39 3.28
VTimeLLM (Huang et al., 2024a) 2.78 3.10 3.40 2.49 2.47 2.85
LITA (Huang et al., 2024b) 2.94 2.98 3.43 2.68 3.19 3.04
Grounded-VideoLLM 3.34 2.94 3.66 3.12 3.14 3.24